# RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives

### v3

@@ 5 Conclusion
Our proposed RACCooN framework newly introduces an auto-descriptive video-to-paragraph-to-video generative framework. RACCooN automatically generates video descriptions by leveraging a multi-granular spatiotemporal pooling strategy, enhancing the model’s ability to recognize detailed, localized video information. RACCooN then uses these enriched descriptions to edit and generate video content, offering users the flexibility to modify content through textual updates, thus eliminating the need for detailed video annotations. These video editing and generation abilities of RACCooN framework highlight notable effectiveness and enable a broader range of users to engage in video creation and editing tasks without the faithfully written textual prompts.
Ethics Statement
The performance of our proposed framework in paragraph generation, video generation, and editing is influenced by the employed pre-trained backbones, including an LLM (Touvron et al., 2023). LLM-empowered video description and photorealistic video creation/editing inherit biases from their training data, leading to several potentially negative impacts, including societal stereotypes, biased interpretation of actions, and privacy concerns. To mitigate these potential negative impacts, it is essential to carefully develop and implement generative and video description models, such as considering diversifying training datasets, implementing fairness and bias evaluation metrics, and engaging communities to understand and address their concerns.
@@ Acknowledgement
We thank Jaemin Cho, Han Lin, Jialu Li, and Elias Stengel-Eskin for the thoughtful discussion. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, and Accelerate Foundation Models Research program. The views contained in this article are those of the authors and not of the funding agency.
@@
  • More details about VPLM dataset collection (Sec. A.1), experimental setups (Sec. A.2), more implementation details (Sec. A.3).
  • Limitations and Negative Societal Impact of RACCooN (Sec. B).
  • Additional analysis including ablations (Sec. C.2, Sec. C.1).
  • Additional qualitative examples with RACCooN on video content editing (Sec. C.3).
@@ Appendix A Experimental Setup
As we mention in Sec. 3.3, to facilitate our model training, we start from open-source video inpainting data (Wu et al., 2024) [^3] MIT License: https://choosealicense.com/licenses/mit/ to build a high-quality dataset that includes the well-structured, detailed caption for both video and each object in the video. Specifically, we leverage GPT-4V 444version 1106 to annotate each video. As shown in Fig. 6, we first convert a video to a superimage (Fan et al., 2021) by concatenating uniformly sampled frames, and we also draw frame IDs on each frame as a visual prompt to present temporal order. Then we prompt the GPT-4V by providing a detailed prompt with in-context examples (left of Fig. 6). In this case, we obtained well-structured, detailed captions that contain both holistic video and local objects (bottom right of Fig. 6). To ensure the annotation quality, we sampled 1 annotated video from each 100 batches and then did a human cross-check, and refined the batch annotations according to the sampled example. Through this pipeline, we obtained 7.2K high-quality quality well-structured, detailed video descriptions with an average of 238.0 words for each video.
Next, to obtain paired object-mask-description triplets for video inpainting model training, we build an automatic detailed object caption and object name matching pipeline using GPT4. As in our base dataset (Wu et al., 2024), we already have class labels for each object mask, we framed this matching as a multi-choice QA to ask GPT4 which object caption can in Fig. 6 matched to the given object classes. We further filtered out the triplets with too small masks ($<1\%$ mask areas.) In this case, we obtained 5.5K object-mask-description triplets with an average of 37.2 words for each object to support our video diffusion model training.
@@ A.2 Benchmarks and Datasets Details
As mentioned in the main paper (Sec. 4.1), we evaluate our proposed RACCooN on various tasks. For video-to-paragraph generation, we test our model on the standard video caption dataset YouCook2 (Zhou et al., 2018a) (validation set) as well as our VPLM dataset. We next test video content editing with three subtasks on our VPLM dataset. Regarding the experiments of incorporating RACCooN with other conditional video generation models, we test RACCooN on diverse videos from ActivityNet, YouCook2, and UCF101. We uniformly selected 100 videos from those 3 datasets to build the test bed. For the experiments of incorporating RACCooN with other video editing models, we follow the previous work (Geyer et al., 2023), and select 30 unique videos from the DAVIS dataset. For each video, we annotate two different types of editing, attribute editing and instance editing. It leads to 60 text-video pairs in our video editing evaluation. We choose object captions that contain the same keywords for editing in human captions to represent the model-generated caption.
@@ A.3 Implementation Details
Metrics: We provide more details about our metrics. CLIP-Text measures the similarity between the edit prompt and the embedding of each frame in the edited video. CLIP-Frame computes the average CLIP similarity between the image embeddings of consecutive frames to measure the temporal coherency. SSIM measures the structural similarity between the original and edited video for evaluating localized editing. $Q_{edit}=CLIP-T/Wrap-Err$, it is a comprehensive score for video editing quality, where $Warp-Err$ calculates the pixel-level difference by warping the edited video frames according to the estimated optical flow of the source video, extracted by FlowNet2.0 (Ilg et al., 2017). For layout planning, we compute the FVD and CLIP-Image similarity between the ground truth and the predicted bounding box.
More Details about Multi-granular Spatiotemporal Pooling. As mentioned Sec. 3.1, we proposed a novel Multi-granular Spatiotemporal Pooling (MGS Pooling) to address the lack of complex spatial-temporal modeling in video. We further provide a more intuitive visualization for our proposed MGS Pooling in Fig. 3. We first adopt a lightweight 165 superpixel predictor that generates superpixels across video frames, then use the overlapping k-means clustering for the obtained video superpixels. In this case, we gather informative cues about various objects and actions for LLM.
@@
Off-shelf Video Editing Models. We utilize TokenFlow (Geyer et al., 2023) and FateZero (Qi et al., 2023) as our video editing tools. TokenFlow generates a high-quality video corresponding to the target text, while preserving the spatial layout and motion of the input video. For SSIM computation, we compute SSIM for region-of-no-interest since we want to keep those regions unchanged. we first mask out regions of interest with the ground truth mask provided by the DAVIS dataset, the we compute SSIM on masked images and conduct mean pooling as the video-level metrics.
Off-shelf Conditional Video Generation Models. We leverage both VideoCrafter (Chen et al., 2023) and DynamiCrafter (Xing et al., 2023) as our video generation backbone. VideoCrafter is one of the SoTA video generation models that can handle different input conditions (image, text). DynamiCrafter is based on the open-source VideoCrafter and T2I Latent Diffusion model (Rombach et al., 2022), and was trained on WebVid10M (Bain et al., 2021), it provides better dynamic and stronger coherence. We adopt VideoCrafter-512 and DynamiCrafter-512 variants. For each video, we use CLIP similarity to retrieve multiple keyframes corresponding to each caption. Those keyframes result in multiple generated video clips via the video generation model. For FVD computation, we conduct mean pooling over those clips to represent a video. We use $k=25$ and $v=6$ for generated captions in all experiments. The experiments are conducted on the 4 × 48GB A6000 GPUs machine.
Appendix B Limitations and Broader Impact
Our proposed RACCooN framework has shown a remarkable ability to interpret input videos, producing well-structured and detailed descriptions that outperform strong video captioning baselines and even ground truths. However, it has the potential to produce inaccuracies or hallucination (Liu et al., 2023a, Wang et al., 2024a, Zhou et al., 2024, Ma et al., 2023) in the generated text outputs. In addition, the performance of our proposed framework in paragraph generation, video generation, and editing is influenced by the employed pre-trained backbones, including an LLM (Touvron et al., 2023), base Inpainting Model (Rombach et al., 2022), Video Diffusion Model (Xing et al., 2023), and Video Editor (Geyer et al., 2023). However, our key contributions are independent of these backbones, and we emphasize that RACCooN’s capabilities can be further enhanced with future advancements in these generative model backbones.
Appendix C Additional Analysis
C.1 Ablation study
The effect of $k$ and $v$ As shown in Tab. 7, we did initialized hyperparameter probing experiments on ActivityNet-Cap and YouCook2 datasets. we observe that all variants of our approach with varying $k$ and $v$ generally achieve improved performance compared to baselines in terms of multiple video captioning metrics: SPICE, BLEU-4, METEOR, and ROUGE. This result demonstrates the efficacy of our multi-granular spatiotemporal pooling approach with a fine to coarse search of video contexts based on superpixels. In addition, we observe that RACCooN framework shows a small gap between variants in each dataset, highlighting the robustness of our approach to the hyperparameter setups and datasets.
@@ C.1 Ablation study
The effect of Superpixel Overlap We introduce overlapping k-means clustering to aggregate video superpixels, capturing a variety of visual contexts while allowing for partial spatiotemporal overlap. To investigate the effect of our suggested overlapping approach, we also evaluate the variant of our framework without overlap (i.e., $v=1$) on video-to-paragraph generation tasks in Tab. 7. As shown, our approach with overlap (i.e., $v>1$) surpasses the non-overlapping version of RACCooN across various scales of visual contexts $k$, as indicated by the video captioning metrics we evaluated. This emphasizes the advantage of permitting overlap in understanding video contexts, which enhances the input video’s comprehension by allowing for diverse and fluent interpretations of local visual regions with surroundings associated at the same time.
@@
C.2 Comparison with Pre-trained Grounding Models
@@
We further investigate the applicability of recent powerful pre-trained visual grounding models (Kirillov et al., 2023, Cheng et al., 2023, Ren et al., 2024). Segment-Anything (Kirillov et al., 2023) and Grounded SAM (Ren et al., 2024) are strong open-ended object segmentation models for images, and we directly compute our localized granular tokens based on their segmentation masks. We select $25$ segmentation masks in total, from uniformly sampled frames in each video for fair comparison with RACCooN ($k=25$). As shown in Tab. 8, these variants of RACCooN achieve improved performance against the best-performing baseline, PG-VL, but are often suboptimal since they focus on regional information and cannot contain temporal information of the videos. Unlike these image-based segmentation methods, SAM-Track (Cheng et al., 2023) generates coherent masks of observed objects over successive frames in videos, by adopting multiple additional pre-trained modules, including GroundingDino (Liu et al., 2023d) and AOT (Yang et al., 2021, Yang & Yang, 2022). We adopt SAM-Track to initialize superpixels in videos and conduct overlapping k-means clustering ($k=25$). Here, we observe that RACCooN with SAM-Track superpixel initialization achieves reasonable performance, and is beneficial for video editing tasks. It enables the model to coherently edit targeted regions in videos with edited keywords.
C.3 Additional Visualizations
@@
Ablation Study Visualization. We illustrate extra visualizations for replacing orecle mask with grounding&tracking tools generated ones for video object removal ( Fig. 17 and  Fig. 18) and changing ( Fig. 19 and  Fig. 20), as well as replacing oracle object boxes with our model-predicted one ( Fig. 21 and  Fig. 22). Our framework shows robust results with LLM planning and off-shelf segmentation tools. We further show the failure cases of removing and changing objects mainly come from the missing mask prediction of the video segmentation masks.

### v4

@@ 5 Conclusion
We newly introduce an auto-descriptive video-to-paragraph-to-video generative approach. We automatically generate video descriptions by leveraging a multi-granular spatiotemporal pooling strategy, enhancing the model’s ability to recognize detailed, localized video information. Our approach then uses these enriched descriptions to edit and generate video content, offering users the flexibility to modify content through textual updates, thus eliminating the need for detailed video annotations. Our video editing and generation abilities highlight notable effectiveness and enable a broader range of users to engage in video creation and editing tasks without the written textual prompts.
Limitations
Our proposed RACCooN has shown a remarkable ability to interpret input videos, producing well-structured and detailed descriptions that outperform strong video captioning baselines and even ground truths. However, it has the potential to produce inaccuracies or hallucination (Liu et al., 2023a; Wang et al., 2024a; Zhou et al., 2024; Ma et al., 2023) in the generated text outputs. In addition, the performance of RACCooN in paragraph generation, video generation, and editing is influenced by the employed pre-trained backbones, including an LLM (Touvron et al., 2023), base Inpainting Model (Rombach et al., 2022), Video Diffusion Model (Xing et al., 2023), and Video Editor (Geyer et al., 2023). However, our key contributions are independent of these backbones, and we emphasize that RACCooN’s capabilities can be further enhanced with future advancements in these generative model backbones.
@@ Acknowledgement
We thank the reviewers and Jaemin Cho, Han Lin, Jialu Li, and Elias Stengel-Eskin for the thoughtful discussion. This work was supported by DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, and Accelerate Foundation Models Research program. The views contained in this article are those of the authors and not of the funding agency.
@@
  • More details about VPLM dataset collection (Sec.˜A.1), experimental setups (Sec.˜A.2), more implementation details (Sec.˜A.3).
  • Additional analysis including ablations (Sec.˜B.2, Sec.˜B.1).
  • Additional qualitative examples with RACCooN on video content editing (Sec.˜B.3).
@@ Appendix A Experimental Setup
Figure 5: Qualitative V2P example of our RACCooN on Sora video.
@@
As we mention in Sec.˜3.3, to facilitate our model training, we start from open-source video inpainting data (Wu et al., 2024) [^3] MIT License: https://choosealicense.com/licenses/mit/ to build a high-quality dataset that includes the well-structured, detailed caption for both video and each object in the video. Specifically, we leverage GPT-4V 444version 1106 to annotate each video. As shown in Fig.˜6, we first convert a video to a superimage (Fan et al., 2021) by concatenating uniformly sampled frames, and we also draw frame IDs on each frame as a visual prompt to present temporal order. Then we prompt the GPT-4V by providing a detailed prompt with in-context examples (left of Fig.˜6). In this case, we obtained well-structured, detailed captions that contain both holistic video and local objects (bottom right of Fig.˜6). To ensure the annotation quality, we sampled 1 annotated video from each 100 batches and then did a human cross-check, and refined the batch annotations according to the sampled example. Through this pipeline, we obtained 7.2K high-quality quality well-structured, detailed video descriptions with an average of 238.0 words for each video.
Next, to obtain paired object-mask-description triplets for video inpainting model training, we build an automatic detailed object caption and object name matching pipeline using GPT4. As in our base dataset (Wu et al., 2024), we already have class labels for each object mask, we framed this matching as a multi-choice QA to ask GPT4 which object caption can in Fig.˜6 matched to the given object classes. We further filtered out the triplets with too small masks ($1\%$ mask areas.) In this case, we obtained 5.5K object-mask-description triplets with an average of 37.2 words for each object to support our video diffusion model training.
@@ A.2 Benchmarks and Datasets Details
As mentioned in the main paper (Sec.˜4), we evaluate our proposed RACCooN on various tasks. For video-to-paragraph generation, we test our model on the standard video caption dataset YouCook2 (Zhou et al., 2018a) (validation set) as well as our VPLM dataset. We next test video content editing with three subtasks on our VPLM dataset. Regarding the experiments of incorporating RACCooN with other conditional video generation models, we test RACCooN on diverse videos from ActivityNet, YouCook2, and UCF101. We uniformly selected 100 videos from those 3 datasets to build the test bed. For the experiments of incorporating RACCooN with other video editing models, we follow the previous work (Geyer et al., 2023), and select 30 unique videos from the DAVIS dataset. For each video, we annotate two different types of editing, attribute editing and instance editing. It leads to 60 text-video pairs in our video editing evaluation. We choose object captions that contain the same keywords for editing in human captions to represent the model-generated caption.
@@ A.3 Implementation Details
In V2P generation, we set $k=[20,25]$ and $v=[5,6]$ for superpixel clustering. We use CLIP-L/14@336 (Radford et al., 2021) as the image encoder and Vicuna-1.5 (Zheng et al., 2024) as the LLM. Our P2V model starts from StableDiffusion-2.0-Inpainting (Rombach et al., 2022). We split the VPLM datasets into train and test sets, with the test set containing 50 unique video-paragraph pairs (for V2P) and 180 mask-object-description triples (for P2V). We manually annotate the editing prompts for the object-changing subtask. We quantitatively compare RACCooN and other baselines on the VLPM test set. To focus on generation results rather than grounding ability, we apply the same ground truth masks and captions to all methods for P2V evaluation. See the supplementary material for more details on datasets, metrics, implementations, ablations, and qualitative analysis.
Metrics: We provide more details about our metrics. CLIP-Text measures the similarity between the edit prompt and the embedding of each frame in the edited video. CLIP-Frame computes the average CLIP similarity between the image embeddings of consecutive frames to measure the temporal coherence. SSIM measures the structural similarity between the original and edited video for evaluating localized editing. $Q_{edit}=CLIP-T/Wrap-Err$, it is a comprehensive score for video editing quality, where WarpErrWarp-Err calculates the pixel-level difference by warping the edited video frames according to the estimated optical flow of the source video, extracted by FlowNet2.0 (Ilg et al., 2017). For layout planning, we compute the FVD and CLIP-Image similarity between the ground truth and the predicted bounding box.
More Details about Multi-granular Spatiotemporal Pooling. As mentioned Sec.˜3.1, we proposed a novel Multi-granular Spatiotemporal Pooling (MGS Pooling) to address the lack of complex spatial-temporal modeling in video. We further provide a more intuitive visualization for our proposed MGS Pooling in Fig.˜3. We first adopt a lightweight 165 superpixel predictor that generates superpixels across video frames, then use the overlapping k-means clustering for the obtained video superpixels. In this case, we gather informative cues about various objects and actions for LLM.
@@
Off-shelf Video Editing Models. We utilize TokenFlow (Geyer et al., 2023) and FateZero (Qi et al., 2023) as our video editing tools. TokenFlow generates a high-quality video corresponding to the target text, while preserving the spatial layout and motion of the input video. For SSIM computation, we compute SSIM for the region-of-no-interest since we want to keep those regions unchanged. We first mask out regions of interest with the ground truth mask provided by the DAVIS dataset, then we compute SSIM on masked images and conduct mean pooling as the video-level metrics.
Off-shelf Conditional Video Generation Models. We leverage both VideoCrafter (Chen et al., 2023) and DynamiCrafter (Xing et al., 2023) as our video generation backbone. VideoCrafter is one of the SoTA video generation models that can handle different input conditions (image, text). DynamiCrafter is based on the open-source VideoCrafter and T2I Latent Diffusion model (Rombach et al., 2022), and was trained on WebVid10M (Bain et al., 2021), it provides better dynamics and stronger coherence. We adopt VideoCrafter-512 and DynamiCrafter-512 variants. For each video, we use CLIP similarity to retrieve multiple keyframes corresponding to each caption. Those keyframes result in multiple generated video clips via the video generation model. For FVD computation, we conduct mean pooling over those clips to represent a video. We use $k=25$ and $v=6$ for generated captions in all experiments. The experiments are conducted on the 4 × 48GB A6000 GPUs machine.
Appendix B Additional Analysis
B.1 Ablation study
The effect of $k$ and $v$ As shown in Tab.˜5, we did initialized hyperparameter probing experiments on ActivityNet-Cap and YouCook2 datasets. we observe that all variants of our approach with varying $k$ and $v$ generally achieve improved performance compared to baselines in terms of multiple video captioning metrics: SPICE, BLEU-4, METEOR, and ROUGE. This result demonstrates the efficacy of our multi-granular spatiotemporal pooling approach with a fine-to-coarse search of video contexts based on superpixels. In addition, we observe that RACCooN shows a small gap between variants in each dataset, highlighting the robustness of our approach to the hyperparameter setups and datasets.
@@ B.1 Ablation study
The effect of Superpixel Overlap We introduce overlapping k-means clustering to aggregate video superpixels, capturing a variety of visual contexts while allowing for partial spatiotemporal overlap. To investigate the effect of our suggested overlapping approach, we also evaluate the variant of our approach without overlap (i.e., $v=1$) on video-to-paragraph generation tasks in Tab.˜5. As shown, our approach with overlap (i.e., $v>1$) surpasses the non-overlapping version of RACCooN across various scales of visual contexts $k$, as indicated by the video captioning metrics we evaluated. This emphasizes the advantage of permitting overlap in understanding video contexts, which enhances the input video’s comprehension by allowing for diverse and fluent interpretations of local visual regions with surroundings associated at the same time.
@@
B.2 Comparison with Pre-trained Grounding Models
@@
We further investigate the applicability of recent powerful pre-trained visual grounding models (Kirillov et al., 2023; Cheng et al., 2023; Ren et al., 2024). Segment-Anything (Kirillov et al., 2023) and Grounded SAM (Ren et al., 2024) are strong open-ended object segmentation models for images, and we directly compute our localized granular tokens based on their segmentation masks. We select $25$ segmentation masks in total, from uniformly sampled frames in each video for fair comparison with RACCooN ($k=25$). As shown in Tab.˜6, these variants of RACCooN achieve improved performance against the best-performing baseline, PG-VL, but are often suboptimal since they focus on regional information and cannot contain the temporal information of the videos. Unlike these image-based segmentation methods, SAM-Track (Cheng et al., 2023) generates coherent masks of observed objects over successive frames in videos, by adopting multiple additional pre-trained modules, including GroundingDino (Liu et al., 2023d) and AOT (Yang et al., 2021; Yang and Yang, 2022). We adopt SAM-Track to initialize superpixels in videos and conduct overlapping k-means clustering ($k=25$). Here, we observe that RACCooN with SAM-Track superpixel initialization achieves reasonable performance, and is beneficial for video editing tasks. It enables the model to coherently edit targeted regions in videos with edited keywords.
B.3 Additional Visualizations
@@
Ablation Study Visualization. We illustrate extra visualizations for replacing orecle mask with grounding&tracking tools generated ones for video object removal ( Fig.˜17 and  Fig.˜18) and changing ( Fig.˜19 and  Fig.˜20), as well as replacing oracle object boxes with our model-predicted one ( Fig.˜21 and  Fig.˜22). RACCooN shows robust results with LLM planning and off-shelf segmentation tools. We further show that the failure cases of removing and changing objects mainly come from the missing mask prediction of the video segmentation masks.
@@
Figure 23: Visualization of text-based video editing. The edited words are marked with Red, and the target words are marked with Green. The RACCooN caption is selected from predicted dense captions. We highlight the region of interest with red dashed-line boxes for comparison.
Figure 24: Visualization of conditional video generation with VideoCrafter. Top (a): we compare generation results conditional on different captions and with the same keyframe. Bottom (b): We leverage multiple keyframes retrieved by different captions to generate multi-scene video. We gray out captions that are not used for retrieval, and highlight captions used for keyframe retrieval. We highlight the region of distortion with red dashed-line boxes for detailed comparison.
Figure 25: Visualization of Masks.